Two charges of equal magnitude and similar sign are 8m apart. What is the electric potential at the mid point between them?

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
k*q/4 + k*q/4
=k*q/2 k=9*10^9


Sher mohammad
b.tech IIT delhi
faculty askiitians